During the previous investigation the ppol became very misaligned to the OPO. I checked today and it is still very misaligned.
First I checked some other stuff. SHG and IRMC are ok.
I tried tapping the ppol EOM, which had no effect on the spectrum. Tapping the OPO case causes a little shift but the OPO case seems to be held tightly. Tapping the OPO input mirror holder causes a large shift with a long transient. Considering that the ppol misalignment was a sudden effect I suspect the PZT may have misaligned the input mirror.
I set up measurement of s-pol BAB in transmission of the OPO (p-pol and s-pol are separated by a Faraday isolator PBS just after the OPO dichroic mirror). I used the stick sensor to measure the power of BAB with the power meter setting 8.8 mW range, BW HI (fixed range and BW HI are necessary for sending the power meter signal to the oscilloscope). The BAB and ppol have a very similar transmission spectrum (figure, yellow is ppol, blue is BAB). Since I didn't touch any of the BAB/ppol alignment optics, this means that the s/p-pol IR beams are nominally co-aligned while the OPO is somehow misaligned. I highly doubt it has anything to do with the crystal so as I suspected it seems the input mirror is misaligned. Normally aligning this inpur mirror is a bit of a pain since the OPO is technically a 3 mirror cavity (first you have to align the laser along the crystal axis and then align the input mirror to optimize the mode structure), however, the input mirror is the only thing that has shifted on this path and it looks like the BAB and ppol are still co-aligned, so I think I can realign the input mirror just by trying to recover ppol/BAB mode matching.